### v3

# What metric to optimize for suppressing instability in a Vlasov-Poisson system?

Abstract
Stabilizing plasma dynamics is an important task in green energy generation via fusion. A common strategy is to introduce an external field to prevent the plasma distribution from becoming turbulent. However, finding such external fields efficiently remains an open question, even for simplified models such as the Vlasov-Poisson (VP) system. In this work, we present an integrated method where we first perform an analytical derivation of the VP system’s dispersion relation to construct a high-quality initial guess for the stabilizing field. This analytically-derived field is then used to initialize a PDE-constrained optimization loop that refines the control to a local optimum. Through extensive numerical experiments, we demonstrate that objective functions evaluated only at the target time—when stable plasma is desired—lead to highly non-convex optimization landscapes, making the global minimum difficult to locate. This behavior arises regardless of the choice of loss function (e.g., KL divergence or electric energy). In contrast, integrating the loss function over time yields a landscape with a convex basin near the global minimum, facilitating convergence. Furthermore, when using electric energy as the objective, the landscape outside this basin is dominated by flat, unphysical local minima, highlighting the critical importance of our analytical approach for generating an initial guess that lies within the global minimum’s basin of attraction.
1 Introduction
Plasma, an ionized gas composed of electrons, ions, and neutral particles, is often referred to as the fourth state of matter [10, 23, 24]. One of its most compelling applications is the generation of clean energy through nuclear fusion [30, 37]. A principal obstacle to realizing practical fusion energy is the inherent instability of plasma dynamics: in many configurations, small perturbations to an equilibrium state can grow exponentially, disrupting the system and rendering a fusion device inoperable. Consequently, developing robust strategies to suppress these instabilities is a critical area of research.
These instabilities are typically mitigated using control strategies, wherein an external actuator—such as a magnetic or electric field—is applied to the plasma to counteract the growth of disruptive modes. Control strategies can be broadly classified by their temporal nature. Dynamic control involves real-time feedback, where the actuator responds continuously to the evolving state of the plasma. This paradigm has been the subject of extensive research, employing both traditional optimal control techniques [1, 2, 7, 18] and modern reinforcement learning (RL) methods [16, 46]. However, dynamic approaches face a fundamental limitation: a timescale mismatch between the rapid growth of certain instabilities and the slower response time of physical actuators [38]. In response to this challenge, static control has recently regained significant interest from the community. This paradigm involves designing a time-independent control to dampen a small initial perturbation, with the goal of stabilizing the system at an ulterior time $T$ [19, 35, 36]. This work focuses on the design of such static controls.
A common approach for designing a static control is to impose an external electric field, $H(x)$, to guide the plasma towards a desired stable distribution [19]. This task can be formulated as a PDE-constrained optimization problem:
$\min_{H}$ $\quad\mathcal{J}(f[H])$ (1.1)
s.t. $\quad\left\{\begin{array}[]{l}\partial_{t}f+v\cdot\nabla_{x}f-(E_{f}+H)\cdot\nabla_{v}f=0\,,\\ E_{f}=\nabla_{x}V_{f}\,,\\ \Delta_{x}V_{f}=1-\rho_{f}=1-\int f\,\mathrm{d}v\,.\end{array}\right$.
Here, the constraint is the Vlasov-Poisson (VP) system, a fundamental kinetic model for collisionless plasma. The function $f:(t,x,v)\in\mathbb{R}_{+}\times\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}_{+}$ is the time-dependent particle distribution in phase space. The system’s evolution is driven by the total electric field, which is a sum of the externally applied control field $H(x)$ and the self-generated electric field $E_{f}(t,x)$. This self-generated field is determined self-consistently: the charge density $\rho_{f}(t,x)=\int f(t,x,v)\,\mathrm{d}v$ determines the electric potential $V_{f}(t,x)$ via the Poisson equation, and $E_{f}$ is determined by its gradient. The objective function $\mathcal{J}$ is a functional designed to quantify the degree of instability in the plasma distribution $f$.
This PDE-constrained optimization framework is a cornerstone of modern control theory and has been employed in numerous studies of plasma stability, both theoretical [18, 26, 27, 35, 34] and numerical [19, 1, 31, 4, 6, 7, 49]. Regardless of the specific physical system, the underlying optimization problem is invariably challenging. The nonlinear nature of the governing PDEs typically leads to highly non-convex objective landscapes, populated with numerous local minima. As such, the careful design of the objective function $\mathcal{J}$ is critical. Different choices for $\mathcal{J}$ emphasize different features of instability [29, 39]; for instance, one metric might penalize the kinetic energy of unstable wave modes, while another might measure the deviation of the entire distribution in phase space from its equilibrium state. This choice directly shapes the optimization landscape and determines which properties of the solution are promoted. This leads to the central question that this manuscript seeks to answer:
What is a good metric for suppressing instabilities in the Vlasov-Poisson system?
In the context of fusion energy, the ultimate goal is to maintain the plasma’s energy in kinetic form, which implies minimizing the energy stored in the self-generated electric field [25, 37]. A common approach in the literature is to target this goal indirectly by minimizing the $L^{2}$-misfit between the final state of the distribution, $f_{T}:=f(T,\cdot,\cdot)$, and the desired equilibrium, $f_{\text{eq}}$ [35, 36]. The rationale behind this approach lies in the fact that if $f_{T}$ perfectly matches the spatially uniform equilibrium $f_{\text{eq}}$, the self-generated electric energy must be zero. However, as demonstrated numerically in [19], this $L^{2}$ objective function is highly non-convex and oscillatory, creating a landscape with a myriad of local minima that can trap optimization algorithms.
This motivates our search for alternative objective functions, which we explore mainly along two axes: the type of metric and its temporal structure. In the first axis, we compare two types of metrics:
  • To measure the deviation from equilibrium, we study using the Kullback-Leibler (KL) divergence as a metric. As $f$ is a probability distribution, the KL-divergence is a more natural information-theoretic measure of disparity than the $L^{2}$ norm.
  • To directly target the underlying physics, we propose minimizing the self-generated electric energy, a macroscopic quantity that is computationally efficient to evaluate and manipulate.
In the second axis, we investigate a crucial structural modification: we compare objective functions evaluated only at the final time $T$ against functions that are integrated over the entire time interval. This allows to test whether incorporating the full dynamics of the system can regularize the problem and appease the non-convexity of the optimization landscape.
However, our numerical landscape analysis reveals that even time-integrated objective functions, while possessing more favorable local properties, remain highly non-convex far from the global minimum. This implies that the success of any gradient-based optimization is critically dependent on the quality of the initial guess for the control field $H$.
We address this issue by analytically building a high-quality initial guess in which we leverage the rich analytical theory of the Vlasov-Poisson system [18]: we consider the linearized regime of the system around the equilibrium state and perform a stability analysis based on the dispersion relation. By identifying the fastest-growing unstable modes (via Fourier and Laplace analysis), we analytically derive a control field specifically tailored to suppress them. While this linear analysis is valid only for short times before non-linear effects dominate, we show numerically that the resulting control provides an initial guess that lies within the basin of attraction of the global minimum for the full non-linear problem (1.1).
In summary, this paper presents an integrated analytical-numerical methodology for suppresing instabilities in the Vlasov-Poisson system. Our numerical simulations, which systematically explore different objective functions, optimization solvers, and initialization strategies, lead to three principal findings:
  • First, we demonstrate that objective functions that integrate information over time yield significantly smoother and more convex-like optimization landscapes around the global minimum compared to those evaluated only at a terminal time.
  • Second, we find that macroscopic metrics such as the electric energy, while physically intuitive, can create spurious, unphysical local minima that trap optimizers, particularly adaptive-steps solvers such as line-search-based methods [41].
  • Third, we show that linear dispersion relation analysis provides a reliable and effective method for constructing an initial guess for the control, capable of placing the optimization within the global minimum’s basin of attraction.
This combined approach allows for the efficient and robust discovery of external fields that effectively suppress the instabilites.
1.1 Related work
Due to its significance, plasma control has attracted considerable research interest, leading to the development of a wide range of control strategies. As one might expect, these strategies are fundamentally shaped by the mathematical models used to describe plasma behavior. These models span a spectrum from macroscopic fluid descriptions—such as magnetohydrodynamics (MHD)[6, 7, 49]—to microscopic kinetic models like the Vlasov–Poisson system. Owing to the high mathematical and computational complexity of kinetic models, many approaches rely on simplifications such as dimensionality reduction[19] or linearization [39, 28, 18] to make the resulting control problems more tractable. These approximations allow the use of optimal control techniques within available computational resources.
Once a model is chosen, another key consideration is the nature of the control strategy itself. Broadly speaking, plasma control can be categorized into static and dynamic control. In static control—such as the approach pursued in this work—the system starts near equilibrium, with only a small perturbation that the control aims to dampen by a prescribed target time [19, 35, 36]. By contrast, dynamic control involves real-time feedback mechanisms that respond continuously to the evolving state of the plasma. Theoretical developments in this direction can be found in [18], while recent experimental implementations have emerged under the optimal control paradigm [1, 2, 7] and the reinforcement learning (RL) paradigm [16, 46]. These RL-based methods recast the control problem in a data-driven framework, training neural networks to predict instabilities [47] and develop reactive strategies. While RL offers the promise of real-time, adaptive control of nonlinear plasma dynamics, it also faces limitations—particularly the mismatch between actuation and measurement timescales (on the order of 25 ms [38]) and the much faster evolution of certain instabilities, which may occur on sub-millisecond or even nanosecond timescales.
Regardless of the control type, implementation in experimental or computational settings typically involves actuating either a magnetic field or an electric field. Magnetic field control is the dominant paradigm in magnetic confinement fusion experiments, particularly for managing MHD-scale behavior in tokamaks [6]. It has also been extended to kinetic models, where theoretical and numerical studies have investigated magnetic control within the Vlasov–Poisson framework to confine plasma away from reactor walls and prevent material damage [1, 2, 35, 36]. In contrast, electric field control—although less common in large-scale experimental devices—has proven powerful in theoretical and computational studies, particularly for kinetic models. Applications include beam shaping in particle accelerators [18] and, most relevant to this work, the suppression of kinetic instabilities via PDE-constrained optimization [19]. Moreover, in one-dimensional configurations, electric fields constitute the only physically viable external control mechanism.
Finally, we should stress that different mathematical definitions of “instability” (objective functions) show drastically different landscapes for a given shared physics problem is not a new phenomenon, and is especially pronounced for wave-type problems, see [20, 50, 13]. In the context of plasma control, though the result is expected, we have not found literature that thoroughly studied this phenomenon.
1.2 Organization
The paper is organized as follows. In Section 2 we properly define the PDE-constrained optimization problem that we aim to solve, along the different possible objective functions that we will test. We also present the forward solver that we will use to solve the forward problem and the two canonical examples that we will present in this work. In Section 3 we briefly present the dispersion relation and a linear stability analysis which provides important information on the frequency and scaling that our control $H$ should have. In Section 4 we thoroughly study the different landscapes for different objectives functions. In Section 5 we make substantial numerical simulations with different objectives, methods, and initializations of the parameters.
2 Preliminaries
In practice, plasma for fusion applications is typically confined within toroidal or axisymmetric domains, such as those found in tokamaks or stellarators, which can be approximated to a circular domain. From a mathematical perspective, this corresponds to treating the $y,z$-domain as constant and expressing the 3D problem as a pseudo-1D problem with periodic boundary conditions, where $(x,v)\in[0,L_{x}]\times[-L_{v},L_{v}]$ with periodic boundary conditions in space and homogeneous Dirichlet boundary conditions in the velocity domain. Under this framework, the Vlasov-Poisson system can be simplified, resulting in the following formulation:
$\left\{\begin{array}[]{lc}\partial_{t}f+v\partial_{x}f-(E_{f}+H)\partial_{v}f=0\,,\\ E_{f}=\partial_{x}V_{f}\,,\\ \partial_{xx}V_{f}=1-\rho_{f}=1-\int f\,\mathrm{d}v\,.\end{array}\right$. (2.1)
It is not hard to see that in the Vlasov-Poisson system defined in (1.1) (and (2.1)), any initial data that is spatially independent is an equilibrium state. In other words, if $f(0,x,v)=f_{0}(x,v)=f_{\text{eq}}(v)$, then $f(t,x,v)=f_{\text{eq}}(v)$ is a solution and, $\partial_{x}f=0$ and $E_{f}=\partial_{x}V_{f}=0$.
We also formally define the electric energy in time $\mathcal{E}_{f}(t)$ as
$\mathcal{E}_{f}(t)=\int_{0}^{L_{x}}[E_{f}(t,x)]^{2}\,\mathrm{d}x\,.$ (2.2)
It is well known that equilibrium states which do not satisfy the Penrose condition [43] are classified as unstable. In such cases, even a small perturbation to the equilibrium state can lead to exponential growth of the self-generated electric field $E_{f}$. In this work, we study two canonical examples of such unstable equilibria that violate the Penrose condition [12, 40]: the Two Stream equilibrium (Section 2.3.1) and the Bump-on-Tail equilibrium (Section 2.3.2).
2.1 PDE-constrained optimization problem
In general, it is difficult to suppress the self-generated electric field for every point in time, so, we attempt to suppress the instability of the Vlasov-Poisson system at a given time $T$ by introducing an external electric field $H(x)$. For this, we define $f[H]=f[H](t,x,v)$ as the solution of (2.1) for a given external electric field $H$ and, our self-generated electric field for that system is now defined as $E_{f[H]}$. Therefore, we want to solve the following PDE-constrained optimization problem.
$\min_{H}$ $\quad\mathcal{J}(f[H])$ (2.3)
[…]
where our objective function will be either of the following four [^1] Another straightforward definition of objective function is to use the $L^{2}$ difference between the equilibrium and the solution $\mathcal{J}(f[H])=\frac{1}{2}\|f_{T}[H]-f_{\text{eq}}\|^{2}_{L^{2}(x,v)}=\frac{1}{2}\int_{-L_{v}}^{L_{v}}\int_{0}^{L_{x}}(f[H](T,x,v)-f_{\text{eq}}(v))^{2}\,\mathrm{d}x\,\mathrm{d}v\,$,. Numerically we find this objective function presents very similar landscape as the KL divergence, as will be discussed in Appendix A.
$\mathcal{J}(f[H])=\text{KL}(f_{T}[H]||f_{\text{eq}})=\int_{-L_{v}}^{L_{v}}\int_{0}^{L_{x}}f[H](T,x,v)\log\left(\frac{f[H](T,x,v)}{f_{\text{eq}}(v)}\right)\,\mathrm{d}x\,\mathrm{d}v\,$, (KL)
$\mathcal{J}(f[H])=\mathcal{E}_{f[H]}(T)=\int_{0}^{L_{x}}[E_{f[H]}(T,x)]^{2}\,\mathrm{d}x\,$, ($E$)
$\mathcal{J}(f[H])=\text{KL}(f_{T}[H]||f_{\text{eq}})=\int_{0}^{T}\int_{-L_{v}}^{L_{v}}\int_{0}^{L_{x}}f[H](t,x,v)\log\left(\frac{f[H](t,x,v)}{f_{\text{eq}}(v)}\right)\,\mathrm{d}x\,\mathrm{d}v\mathrm{d}t\,$, (KLT)
$\mathcal{J}(f[H])=\int_{0}^{T}\mathcal{E}_{f[H]}(t)\mathrm{d}t=\int_{0}^{T}\int_{0}^{L_{x}}[E_{f[H]}(t,x)]^{2}\,\mathrm{d}x\,\mathrm{dt}\,$, (EET)
where $E_{f[H]}$ and $\mathcal{E}_{f[H]}$ are as defined in (2.1) and (2.2) respectively. We note that these four objective functions can be grouped into two categories: one based on electric energy and the other based on KL divergence. The former is a natural, physics-inspired choice, while the latter is a valid approach that has been considered in a recent work on control for the Vlasov-Poisson system [8].
To simplify the computations we define the external field $H(x)$ as a linear combination of cosine and sine basis functions:
$H(x;\bm{a},\bm{b})$ $=\sum_{k=1}^{N}a_{k}\cos\left(\frac{2\pi kx}{L_{x}}\right)+b_{k}\sin\left(\frac{2\pi kx}{L_{x}}\right)$ (2.5)
$=\sum_{k=1}^{N}a_{k}\cos\left(kk_{0}x\right)+b_{k}\sin\left(kk_{0}x\right)\,$,
where $\bm{a}=(a_{1},...,a_{N})$ and $\bm{b}=(b_{1},...,b_{N})$ represent the parameters to be determined. To simplify notation we will drop the dependence on $\bm{a}$ and $\bm{b}$ for $H$, and our problem becomes,
$\min_{(\bm{a},\bm{b})}$ $\quad\quad\quad\mathcal{J}(f[H])$ (2.6)
[…]
$\,(\bm{a},\bm{b})\in\mathbb{R}^{N}\times\mathbb{R}^{N}\,.$
2.2 Solving the forward problem
In order to solve (2.6) we must be able to solve (2.1) in an efficient and stable way. In the past, many methods have been developed: Eulerian methods [21], Lagrangian methods [51] and, semi-Lagrangian methods [14, 45]. A general summary can be found in [22]. In this work we will use the latter mainly because they are fully explicit and unconditionally stable (they do not suffer from a CFL condition [5, 11]).
In a nutshell, semi-Lagrangian methods trace the characteristics exactly and perform an interpolation since the translated solution may not coincide with the grid. The literature proposes different ways to perform this interpolation, with spline based [14, 48, 21] and Fourier-based methods [32, 33] being the most famous. More recently, the interpolation has been done following a discontinuous Galerkin approach [15, 44, 45, 17, 19]. In our case, we deploy a simple linear interpolation.
We now describe the semi-Lagrangian method that we implemented. We first perform a time splitting to treat the computations in space and velocity separately such that:
$\partial_{t}f+v\partial_{x}f=0$ (2.7a)
$\partial_{t}f-(E_{f}+H)\partial_{v}f=0$ (2.7b)
One can solve for each of these equations by using the method of characteristics as long as we have an initial condition. For this, we discretize in time such that our time step is $\Delta t$ and $t^{n}:=n\Delta t$ is the $n$-th time step taken. We also discretize the phase-space $(x,v)$ using equispaced grid points such that we have $M_{x}$ and $M_{v}$ points for each. For simplicity, we will use the same amount for each so $M=M_{x}=M_{v}$, and our grid becomes $\{x_{i},v_{j}\}_{i,j=0}^{M-1}$.
Since we want to make sure that we take one time step to solve the whole VP system we first solve for (2.7a) using half a step. In this step, we will do a linear interpolation because we may have to evaluate our solution outside of our grid points, so we get a numerical approximation of the solution which we note by $\mathsf{f}$. In other words, if we have an approximation of the solution at $t^{n}$ and want to get an approximation at $t^{n+\frac{1}{2}}$, we compute
$\mathsf{f}^{\ast}(x_{i},v_{j})=\mathsf{f}(t^{n},x_{i}-v_{j}\Delta t/2,v_{j})\,,\quad\text{for }i,j=0,1,...,M-1\,.$
Next, we use this solution to compute the self-generated electric field $E_{\mathsf{f}^{\ast}}$. For this, we solve
$\left\{\begin{array}[]{rcl}E_{\mathsf{f}^{\ast}}&=&\partial_{x}V_{\mathsf{f}^{\ast}}\\ \partial_{xx}V_{\mathsf{f}^{\ast}}&=&1-\rho_{\mathsf{f}^{\ast}}\end{array}\right.\,$. (2.8)
using a pseudo-spectral method. We apply the Fourier transform so,
$\left\{\begin{array}[]{rcl}\hat{E}_{\mathsf{f}^{\ast}}&=&2\pi i\xi\hat{V}_{\mathsf{f}^{\ast}}\\ 4\pi^{2}\xi^{2}\hat{V}_{\mathsf{f}^{\ast}}&=&\hat{\rho}_{\mathsf{f}^{\ast}}\,,\quad(\forall\xi\neq 0)\end{array}\right.\,$.
Since we can compute $\hat{\rho}_{\mathsf{f}^{\ast}}$ by knowing $\mathsf{f}^{\ast}$ from the previous step, we have that
$\hat{E}_{\mathsf{f}^{\ast}}(t,\xi)=-\frac{1}{2\pi i\xi}\hat{\rho}_{\mathsf{f}^{\ast}}(t,\xi)\,.$
Then, applying the inverse Fourier transform gives $E_{\mathsf{f}^{\ast}}$. With this, we can solve for (2.7b) in the same way we solved for (2.7a) for the whole time step so,
$\mathsf{f}^{\ast\ast}(x_{i},v_{j})=\mathsf{f}^{\ast}(x_{i},v_{j}+(E_{\mathsf{f}^{\ast}}+H)\Delta t)\,,\quad\text{for }i,j=0,1,...,M-1\,.$
Finally, we take the remaining half of the time step on the space variable so
$\mathsf{f}(t^{n+1},x_{i},v_{j})=\mathsf{f}^{\ast\ast}(x_{i}-v_{j}\Delta t/2,v_{j})\,,\quad\text{for }i,j=0,1,...,M-1\,.$
We summarize the semi-Lagrangian method described above in Algorithm 1222https://github.com/maguerrap/Vlasov-Poisson.
Algorithm 1 Semi-Lagrangian with interpolation
[…]
3:for n=0,1,,Nn=0,1,...,N do
4:  Compute $\mathsf{f}^{\ast}(x_{i},v_{j})=\mathsf{f}(t^{n},x_{i}-v_{j}\Delta t/2,v_{j})$, for i,j=0,1,,M1i,j=0,1,...,M-1.
5:  Compute $E_{\mathsf{f}^{\ast}}$ by solving (2.8).
6:  Compute $\mathsf{f}^{\ast\ast}(x_{i},v_{j}+(E_{\mathsf{f}^{\ast}}+H)\Delta t)$, for i,j=0,1,,M1i,j=0,1,...,M-1.
7:  Compute $\mathsf{f}(t^{n+1},x_{i},v_{j})=\mathsf{f}^{\ast\ast}(x_{i}-v_{j}\Delta t/2,v_{j})$, for i,j=0,1,,M1i,j=0,1,...,M-1.
8:end for
2.3 Two canonical examples
Plasma system, or the system modeled by the Vlasov-Poisson equation set, can have many different kinds of instabilities. We only study two most prominent ones: Two Stream instability and Bump-on-Tail instability. Both instabilities violate the Penrose condition and are regarded as two canonical examples.
2.3.1 Two Stream example
As a first example, we consider the Two Stream equilibrium. This distribution arises when two streams (or beams) of charged particles moving relative to each other are interpenetrating, causing instability [3]. Mathematically, the equilibrium distribution is given by
$f_{\text{eq}}(v)=\frac{\alpha\exp\left(-\frac{1}{2}(v-\mu)^{2}\right)+(1-\alpha)\exp\left(-\frac{1}{2}(v+\mu)^{2}\right)}{\sqrt{2\pi}}\,.$ (2.9)
It is a distribution composed of two Gaussians having opposite centers. In our simulation we set $\alpha=0.5$ and $\mu=2.4$.
The initial data is set to be a small perturbation from this equilibrium:
$f_{\text{iv}}(x,v)=\left(1+\varepsilon\cos(k_{0}x)\right)f_{\text{eq}}(v)\doteq f_{\text{eq}}(v)+f_{\text{p}}(t=0,x,v)\,$, (2.10)
with the initial perturbation $f_{\text{p}}(t=0,x,v)$ being of order $\varepsilon=0.001$.
In Figure 1 we plot our simulation (2.1) using Algorithm 1 without an external electric field (i.e., $H\equiv 0$). We simulate until $T=30$ – a time one can observe strong instability starting to appear. The simulation is done over a $L_{x}=10\pi$ and $L_{v}=6.0$ box.
Figure 1: Simulation of (2.1) with $H\equiv 0$ for the Two Stream equilibrium. From left to right we have $f_{\text{eq}}(v)$, $f(T=30,x,v)$, $E_{f}(t,x)$ and $\mathcal{E}_{f}(t)$.
2.3.2 Bump-on-Tail example
The Bump-on-Tail equilibrium is another steady state that is heavily investigated. This distribution appears when a small group of fast-moving electrons (a “bump”) is superimposed on the tail of the background electron velocity distribution and this can cause instability [42]. This can happen due to particle injection or wave-particle interactions. The equilibrium is defined as
$f_{\text{eq}}(v)=\frac{9}{10\sqrt{2\pi}}\exp\left(-\frac{1}{2}(v-\bar{v}_{1})^{2}\right)+\frac{\sqrt{2}}{10\sqrt{\pi}}\exp\left(-2(v-\bar{v}_{2})^{2}\right)\,.$ (2.11)
with the equilibrium parameters set to
$\bar{v}_{1}$ $=-3.0,\quad\bar{v}_{2}=4.5.$
This equilibrium is also not stable. We add an initial perturbation to it as the initial data:
$f_{\text{iv}}(x,v)=f_{\text{eq}}(v)+\underbrace{\frac{\sqrt{2}\,\varepsilon}{10\sqrt{\pi}}\exp\left(-2(v-\bar{v}_{2})^{2}\right)\cos(0.1x)}_{\doteq f_{\text{p}}(t=0,x,v)}\,$, (2.12)
with the initial perturbation $f_{\text{p}}(t=0,x,v)$ being of order $\varepsilon=0.001$, the electric energy will exponential grow in time. In Figure 2 we simulate (2.1) using Algorithm 1 without an external electric field (i.e., $H\equiv 0$). The solution demonstrates strong electric energy growth at about $T=40$ when we terminate the simulation. The simulation is conducted over a $L_{x}\times L_{v}=20\pi\times 9.0$ box.
Figure 2: Simulation of (2.1) with $H\equiv 0$ for the Bump-on-Tail equilibrium. From left to right we have $f_{\text{eq}}(v)$, $f(T=40,x,v)$, $E_{f}(t,x)$ and $\mathcal{E}_{f}(t)$.
3 Dispersion relation and linear stability analysis
Examining the dispersion relation is a classical approach to investigating the stability properties of dynamical systems [39]. Mathematically, this involves linearizing the system around a desired equilibrium state and analyzing the spectrum of the resulting linear PDE operator. If the operator has spectrum with positive real parts, the system exhibits exponential divergence from equilibrium. To counteract this, we introduce a control term $H$ aimed at modifying the spectrum—specifically, to eliminate its unstable (positive) components and thereby keep the system close to the desired state.
This strategy was proposed and studied in depth in [18], where the dispersion relation was derived both with and without an external field, and a general procedure was outlined to design a field capable of suppressing growth. However, the method targets complete suppression by eliminating all unstable roots, which is a stringent requirement. Achieving this demands highly flexible external fields—indeed, the designs in [18] result in fields that oscillate significantly in time. While theoretically sound, such time-dependent fields are physically impractical and incompatible with our assumption (2.5).
In contrast, we seek to impose a far less flexible external field: one that contains only finitely many Fourier modes and is time-independent. This constraint necessitates a relaxation of the original goal. Rather than eliminating all unstable roots, we shift our focus to suppressing the modes that grow the fastest. Our approach begins by identifying the specific modes that would exhibit the strongest exponential growth in the absence of control, then targeting their associated roots for elimination. This analysis builds on our previous derivation of the dispersion relation, now adapted to account for time-independent control.
Ultimately, the system remains nonlinear, so previously unexcited modes may be triggered over time. As a result, the present strategy is effective only in the short term. For long-term suppression, a return to a full PDE-constrained optimization framework is necessary. Nonetheless, our analysis offers a good initial guess for such optimization-based methods.
In the following section, we begin by reviewing the dispersion relation derived in [18], as presented in subsection 3.1. We then extend the analysis to the case of a time-independent external field, and propose our strategy in subsection 3.2. It provides a good initial guess for further optimization.
3.1 Dispersion relation and control strategy when $H=H(t,x)$
We firstly briefly review the dispersion relation for the Vlasov–Poisson system subject to an external field $H(t,x)$ that varies in both time and space. The original derivation is complete in [18].
To do so, we first linearize the system around the desired equilibrium state, writing $f=f_{\text{eq}}+f_{\text{p}}$, then omitting the higher order terms in the expansion, the leading order solution to $f_{\text{p}}$ is:
$\left\{\begin{aligned} &\partial_{t}f_{\text{p}}(t,x,v)+v\partial_{x}f_{\text{p}}(t,x,v)-\left[E_{f_{\text{p}}}(t,x)+H(t,x)\right]\partial_{v}f_{\text{eq}}(v)=0\,,\\ &E_{f_{\text{p}}}(t,x)=\partial_{x}V_{f_{\text{p}}}(t,x)\,,\\ &\partial_{xx}V_{f_{\text{p}}}(t,x)=-\rho_{\text{p}}(t,x)=-\int f_{\text{p}}(t,x,v)\,\mathrm{d}v\,.\end{aligned}\right$. (3.1)
Three crucial transforms will be heavily leaned on:
  • Fourier transform in space, e.g.
    $\hat{F}(t,k)=\int_{-\infty}^{\infty}F(t,x)e^{-ikx}\,\mathrm{d}x\,.$
  • Fourier transform in space and velocity, e.g:
    $\hat{F}(t,k,m)=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}F(t,x,v)e^{-ikx}e^{-imv}\,\mathrm{d}x\,\mathrm{d}v\,.$
  • Laplace transform in time:
    $L[F](s)=\int_{0}^{\infty}e^{-st}F(t)\,\mathrm{d}t\,.$ (3.2)
Repeatedly deploying these transforms, one can deduce an important identity:
$\left(L[\hat{\rho}_{\text{p}}(\cdot,k)](s)-L[\hat{S}(\cdot,k)](s)\right)\left(1+L[\hat{U}(\cdot,k)](s)\right)=L[\hat{U}(\cdot,k)](s)\left(-L[\hat{S}(\cdot,k)](s)+ikL[\hat{H}(\cdot,k)](s)\right).$ (3.3)
where
  • $\hat{U}(t,k)$ is defined as:
    $\hat{U}(t,k):=t\hat{f}_{\text{eq}}(kt)\,.$ (3.4)
    It is a function of $t$ and $k$, and is uniquely determined by the equilibrium $f_{\text{eq}}$.
  • $\hat{S}$ is defined as
    $\hat{S}(t,k):=\hat{f}_{\text{p}}(0,k,kt)\,.$ (3.5)
    Or equivalently $S(t,x)=\int f_{\text{p}}(t,x-vt,v)\,\mathrm{d}v$. It is the free-streaming solution.
According to Mellin inversion Laplace transform formula, $\rho_{\text{p}}$ will have exponential growth if $L[\hat{\rho}_{\text{p}}]$ has a singularity on $\Re(s)>0$, while the source term on the right hand side of (3.3) does not vanish at the same roots. As a consequence, to eliminate the exponential growth, one should also force the right hand side to be zero at the roots of $1+L[\hat{U}(\cdot,k)](s)$. In light of this discussion, in [18], the authors propose to set the control $H$ according to:
$L[\hat{S}(\cdot,k)](s)+ikL[\hat{H}(\cdot,k)](s)=h\left(1+L[\hat{U}(\cdot,k)](s)\right),$ (3.6)
where $h$ is a factor so that:
$h:\mathbb{C}\to\mathbb{C}\quad\text{s.t.}\quad h(0)=0\quad\text{and}\quad\lim\limits_{|x|\to 0}\frac{|h(\cdot)|}{|\cdot|}=c\ \ \text{for some}\ c\in\mathbb{C}\,.$ (3.7)
This design effectively eliminates the positive roots, and thus the PDE’s instability is suppressed. Further analytical derivations and comprehensive numerical validations are provided in [18].
3.2 Dispersion relation and control strategy when $H=H(x)$
It is evident that the control $H$ designed above has both time and space dependence. However, the $H$ we are seeking for only has spatial dependence, rendering these analysis results not immediately useful. In this section, we modify the analysis above and re-derive the dispersion relation with $H$ assumed to be time-independent.
Similar to the derivation above, we apply the Fourier–Laplace transform and find:
$\bigl{(}L[\hat{\rho}_{\mathrm{p}}(\cdot,k)](s)-L[\hat{S}(\cdot,k)](s)\bigr{)}\,\bigl{(}1+L[\hat{U}(\cdot,k)](s)\bigr{)}=L[\hat{U}(\cdot,k)](s)\,\bigl{(}-L[\hat{S}(\cdot,k)](s)+\tfrac{ik\,\hat{H}(k)}{s}\bigr{)}.$ (3.8)
In the derivation, we used $L[1](s)=\frac{1}{s}$ and the fact that $H$ does not have time dependence.
Following the strategy as in (3.6), we should set
[…]
for all complex frequencies $s$. The equation is pointwise in $(k,s)$ while the function to be sought for only varies in $k$, rendering no solution that satisfies the entire equation. However, as discussed, in this regime when we have a far less flexible external field, we ought to relax our original goal. In this particular setting, we are to find the fastest growing mode and aim at suppressing that mode only.
Mathematically, to find the fastest growing mode, we set
$s_{0}(k)=\text{argmax}\{\Re(s)>0:\;1+L\bigl{[}\hat{U}(\cdot,k)\bigr{]}(s)=0\}$ (3.9)
for every fixed $k$. This is to identify the Laplace index who are roots for the operator and has the largest real component. We then require $\hat{H}$ only to eliminate that specific component: $L\bigl{[}\hat{S}(\cdot,k)\bigr{]}(s_{0}(k))\;+\;\frac{ik\,\hat{H}(k)}{s_{0}}=0$. This yields
$\hat{H}(k)\;=\;\frac{s_{0}(k)\,L\bigl{[}\hat{S}(\cdot,k)\bigr{]}(s_{0}(k))}{ik}\,.$ (3.10)
We summarize our strategy as follows. For a given equilibrium distribution $U$ and the initial perturbation $S$, we:
  1. 1.
    compute (3.9) for $s_{0}(k)$ for all $k$ seen in the initial perturbation;
  2. 2.
    compute the spectral amplitude $\hat{H}(k)$ using (3.10).
Since $\hat{S}$ is determined, according to (3.5), by the initial perturbation. By definition (2.10) and (2.12), it is of order $\mathcal{O}(\varepsilon)$, hence the guessed control $H$ should also be of $\mathcal{O}(\varepsilon)$.
Example 3.1.
We apply this strategy on Two Stream instability and Bump-on-Tail instability.
  • Two Stream instability: The initial perturbation (2.10) only sees $k=\pm 1$, so we only compute the associated $s_{0}(\pm 1)$. As shown in Figure 3 (left panel) that plots $|1+L(\hat{U}(\cdot,1))|(s)$, we find that the numerical root that has the biggest real component is
    $s_{0}(\pm 1)=0.236\,.$
    Substituting it into (3.10), we have:
    $\hat{H}(1)=-3.15768\times 10^{-9}+0.000643706\,i,\quad\hat{H}(-1)=-3.15768\times 10^{-9}-0.000643706\,i,$
    and thus
    $H(x)=-0.00128741\,\sin(k_{0}x)-6.31537\times 10^{-9}\,\cos(k_{0}x)$
  • Bump-on-tail instability: The initial perturbation (2.12) only sees $k=\pm 1$, so we only compute the associated $s_{0}(\pm 1)$. As shown in Figure 3 (right panel) that plots $|1+L(\hat{U}(\cdot,1))|(s)$, we find that the numerical root that has the biggest real component is
    $s_{0}(\pm 1)=0.230\mp 0.324\,i\,.$
    Substituting it into (3.10), we have:
    $\hat{H}(1)=7.05072\times 10^{-4}+1.60134\times 10^{-4}\,i,\quad\hat{H}(-1)=7.05072\times 10^{-4}-1.60134\times 10^{-4}\,i.$
    and thus
    $H(x)=-0.000320268\,\sin(k_{0}x)+0.00141014\,\cos(k_{0}x)\,.$
Figure 3: Norm of $\|1+\mathcal{L}[U](\cdot,1)(s)\|$: the Two Stream (left) and the Bump-on-Tail (right) examples. In Two Stream, the minimum occurs at $s_{0}=0.236+0i$ with $\|1+\mathcal{L}[U](\cdot,1)(s_{0})\|=9.564044\times 10^{-4}$. In Bump-on-Tail, the minimum occurs at $s_{0}=0.230-0.324i$ with $\|1+\mathcal{L}[U](\cdot,1)(s_{0})\|=3.756278\times 10^{-3}$.
These calculations yield a strong initial guess for the external field $H$. When applied to equation (2.1), we observe significant suppression effects, as illustrated in Figure 4. However, it is important to note that the electric energy eventually exhibits exponential growth over long times. This reflects the onset of nonlinear dynamics, which cannot be fully addressed by our initial design. As such, complete suppression requires the use of a PDE-constrained optimization framework. The field $H$ constructed here serves as an effective initial guess for such optimization, which we will discuss in detail in Section 5.
Figure 4: Simulation of (2.1) for the Two Stream equilibrium up to $T=30$ (top) and Bump-on-Tail equilibrium up to $T=40$ (bottom). In each row, from left to right: $f_{\text{eq}}(v)$, $f(T,x,v)$ (for $H\equiv 0$), $f(T,x,v)$ (for $H$ obtained from Example 3.1) and $\mathcal{E}_{f}(t)$ (for $H\equiv 0$ and $H$ obtained from Example 3.1).
4 Landscape analysis of the objective function
In this section, we examine the landscape of the objective function. This is to evaluate the dependence of the four objective functions (defined in (KL), (EE), (KLT) and (EET)) on parameters of the control (defined in (2.5)). Landscape analysis requires a sweep of PDE solves over all parameter choices in a pre-defined domain, and thus is computationally demanding.
In Section 4.1 and Section 4.2 respectively, we provide the landscape for choices of two different parameter settings for the two different canonical examples (Sections 2.3.1 and 2.3.2) respectively.
4.1 Two Stream example
We report results for the Two Stream instability problem in this section. Here we have the following two choices of parameter settings:
  • Choice A:
    $H(x)=b_{1}\sin(k_{0}x)\,$, (4.1)
  • Choice B:
    $H(x)=b_{1}\sin(k_{0}x)+b_{2}\sin(2k_{0}x)\,.$ (4.2)
For Choice A, with one parameter, we conduct a sweeping of PDE solutions by setting $b_{1}$ to be in the pre-defined range of [0.1,0.1][-0.1,0.1], and compute the evaluation of the four objective functions ((KL), (EE), (KLT) and (EET)). The result is plotted in Figure 5. Evidently, in this 1D situation, both (EE) and (KL) that only see the last frame of the solution give non-convex behavior in the neighborhood of the minimum, whereas (EET) and (KLT) that see the whole solution in time demonstrates a convex profile. This strongly suggests the information in-time is crucial in determining the convexity of the landscape of the objective functions.
Figure 5: Landscape of the Two Stream instability on the domain [0.07,0.07][-0.07,0.07] for $b_{1}$ for (4.1). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
Then, for Choice B with two parameters, we again perform a sweeping of PDE solutions by setting $(b_{1},b_{2})$ in three pre-defined domains:
  • Far field: $1,1]\times[-1,1]$;
  • Mid-range: $0.1,0.1]\times[-0.1,0.1]$;
  • Near field: $0.003,0.001]\times[-0.003,0.001]$.
The wordings (“Far field, Mid-range, Near field”) are selected based on the conclusions drawn from Section 3 of dispersion relation, where it was suggested the optimal $H$ should be in the range of $O(\varepsilon)$. In this case, considering (2.10), $\varepsilon=0.001$. In Figure 6, we plot the values for the four different objective functions in these three domains, and some conclusions are drawn:
  • The landscape close to the global minimum (near-field) is convex for all four objective functions.
  • In the mid-range set, both plots that only use last-frame information ((KL)-(EE)) present a pronounced bulge around the minimum that highlights a clear non-convex behavior along the first mode direction ($b_{1}$ variable), while the plots that integrate in time ((KLT)-(EET)) eliminate these bulge and present a convex objective landscape.
  • In the far field, both (KL) and (EE) are highly oscillatory, making the global minima hard to find. In comparison, the objective functions that integrates in time are much smoother. Comparing KL and electric energy, we also see that, in both final-frame and integrated-in-time case, objective functions defined by electric energy present minima for large values of $b_{2}$, suggesting large external electric field can also suppress instability. However, we should point out that this is a non-physical setting: The imposed electric field $H$ is large, and completely dominates the dynamics.
Figure 6: Landscape of the Two Stream instability with on the domain $1,1]^{2}$(top), $0.1,0.1]^{2}$(center) and, $0.003,0.001]^{2}$(bottom) for $(b_{1},b_{2})$ for (4.2). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
4.2 Bump-on-Tail example
In this section, we present results for the Bump-on-Tail instability problem with the following two choices of parameter settings:
  • Choice C:
    $H(x)=a_{1}\cos(k_{0}x)\,$, (4.3)
  • Choice D:
    $H(x)=a_{1}\cos(k_{0}x)+b_{1}\sin(k_{0}x)\,.$ (4.4)
For Choice C, we perform a parameter sweep of the PDE solutions by varying $a_{1}$ over the pre-defined interval [0.1,0.1][-0.1,0.1]. We then evaluate four different objective functions ((KL), (EE), (KLT) and (EET)), in analogous to the Two Stream example. The results are displayed in Figure 7. Notably, in this one-dimensional setting, both (KL) and (EE) exhibit non-convex landscape, whereas (KLT) and (EET) display a convex profile, mirroring the behavior observed in the Two Stream example in Figure 5. This observation resonates with the situation for the Two Stream instability, and suggests information in-time is crucial in convexifying the objective functions.
Figure 7: Landscape of the Bump-on-Tail instability on the domain [0.07,0.07][-0.07,0.07] for $a_{1}$ for (4.3). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
For Choice D, we extend the analysis to a two-parameter sweep by varying the pair $(a_{1},b_{1})$ over three pre-defined domains ($1,1]^{2}$, $0.1,0.1]^{2}$, and $0.001,0.003]^{2}$). The near-field region is shifted due to the new location of the global minimum in this case.
Figure 8 displays the values of the four objective functions ((KL), (EE), (EET) and (KLT)) across the three domains. The conclusions drawn from the Two Stream example continue to hold in the near-field and mid-range cases. In the far-field region, (EET) and (KLT) continue to be more convex than those of (KL) and (EE). Moreover, they do not exhibit global minima for large values of $a_{1}$ and $b_{1}$ as seen in Figure 6 for far-field region. This observation indicates that an excessively strong external electric field does not always produce a correspondingly small self-generated electric field.
Figure 8: Landscape of the Bump-on-Tail instability with on the domain $1,1]^{2}$(top), $0.1,0.1]^{2}$(center) and, $0.001,0.003]^{2}$(bottom) for $(a_{1},b_{1})$ for (4.4). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
5 Solving the PDE-constrained optimization problem
In this section, we present the numerical results obtained from solving (2.6) using Algorithm 1 to address the forward system described by (2.1). The computation was performed on an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory. For the forward simulation, the phase space was discretized with $M=256$ mesh points, and the time domain was discretized using a time step of $\Delta t=0.1$. Each PDE solve can be completed within 5 seconds. Optimization was carried out using automatic differentiation, facilitated by the Python library JAX [9], to compute the required gradients.
The plots shown in Section 4 suggest that the landscape for the objectives (KLT) or (EET) are rather similar. Therefore we expect the performance of optimization solvers with these objectives would be similar as well. The rest of this section, we perform numerical studies only on (KL), (EE) and (EET).
The landscape plots only provide some suggestions on the performance of optimization solvers. Indeed, for local solvers such as gradient descent, that can only see local behavior of objective functions, the solutions are trapped in local minima, and thus placing the initial guess in the global basin is crucial. On the other hand, solvers have their non-convex extensions. One such possibility is to integrate the line-search technique, which allows iterations to jump out of local minima. This can potentially relax the requirement for a good initial guess.
Another strategy typically deployed in optimization is “over-parameterization”. This is to increase the number of unknowns and adjustable parameters, and thus formulating the problem in a higher-dimensional space. In this setting, it is very likely that the dimension of the global optima manifold increases, and so the chance of obtaining optimal points.
For these reasons, we numerically investigate the solvers’ performance on two physics scenarios, Two Stream instability and Bump-on-Tail instability, respectively, in three distinct dimensions of variations:
  • Solvers can be either local or adaptive/global;
    • Local solver. There are many kinds of local solvers. We adopt the simplest gradient descent with constant stepsize. The stepsize is small enough (1e81e-8, 1e91e-9) that one can essentially view the optimization iteration as a gradient flow along the landscape.
    • Adaptive/global solver. There are also many choices of adaptive solvers. To be more comparable with the choice of local solver (GD), we select GD with line-search correction as our adaptive solver. The line-search will satisfy the strong Wolfe condition that we briefly describe below.
      Denoting the GD with stepsize $\alpha_{n}$ at $n$-th iteration for minimizing a function $F$:
      $x_{n+1}=x_{n}-\alpha_{n}\nabla F(x_{n})\,.$
      Line-search is imposed to ensure that the choice of $\alpha_{n}$ provides sufficient decrease. The “sufficiency” is quantified by the following:
      $F(x_{n}-\alpha_{n}\nabla F(x_{n}))$ $\leq F(x_{n})-c_{1}\alpha_{n}\|\nabla F(x_{n})\|^{2}$
      $|[\nabla F(x_{n}-\alpha_{n}\nabla F(x_{n}))]^{\top}\nabla F(x_{n})|$ $\leq c_{2}\|\nabla F(x_{n})\|^{2}$
      for $0<c_{1}<c_{2}<1$. Namely, either the value of the new update is sufficiently smaller than the current value, or the new update has a gradient almost parallel to that of the old one. In particular, we use $c_{1}=10^{-4}$ and $c_{2}=0.9$.
  • Initial guess can be close, or far away from the optimum points. More specifically, the three kinds of initializations are
    • Far initialization: $(\bm{a},\bm{b})\in[-1,1]^{2N}$,
    • Mid-range initialization: $(\bm{a},\bm{b})\in[-0.05,0.05]^{2N}$,
    • Near initialization: $(\bm{a},\bm{b})\in[-0.003,0.001]^{2N}$ for Two Stream and $(\bm{a},\bm{b})\in[-0.001,0.003]^{2N}$ for Bump-on-Tail.
    The scaling are termed “far”, “mid-range” and “near” are closely tied to the landscapes we found on Section 4.
  • Optimization conducted with either $2$ (under or properly parameterized) or $14$ (over-parameterized) variables.
    According to the definition of the control term (2.5), $H$ is composed of Fourier modes. The terminology “under/properly” or “over” have blurred scientific meanings. Roughly speaking, a system is “over-parameterized” if one can find a configuration that already performs optimization well with a smaller set of parameters. We perform the study with
    • Under/properly parameterized:
      $H(x;b_{1},b_{2})=b_{1}\sin(k_{0}x)+b_{2}\sin(2k_{0}x)\,,\quad\text{(Two Stream)}$ (5.1)
      $H(x;a_{1},b_{1})=a_{1}\cos(k_{0}x)+b_{1}\sin(k_{0}x)\,,\quad\text{(Bump-on-Tail)}$ (5.2)
    • Over-parameterized:
      $H(x;\bm{a},\bm{b})=\sum_{k=1}^{14}\big{(}a_{k}\cos\left(kk_{0}x\right)+b_{k}\sin\left(kk_{0}x\right)\big{)}$ (5.3)
      where $(\bm{a},\bm{b})=((a_{1},...,a_{14}),(b_{1},...,b_{14}))$.
5.1 Two Stream example
We summarize results for controlling Two Stream instability, using the equilibrium state and initial condition presented in (2.9)-(2.10). All numerical results are presented in Table 1 that shows the final distribution function $f_{T}[H]$ with $H$ found by the optimization solver. A more detailed summary of each experiment can be found in Appendix B.1. The stepsize used for the constant stepsize gradient descent is $10^{-8}$.
Under-parametrized (5.1) Over-parametrized (5.3)
Init. Step (KL) ($E$) (EET) (KL) ($E$) (EET)
Far Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 10) (Fig. 16) (Fig. 22) (Fig. 28) (Fig. 34) (Fig. 40)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 11) (Fig. 17) (Fig. 23) (Fig. 29) (Fig. 35) (Fig. 41)
Mid Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 12) (Fig. 18) (Fig. 24) (Fig. 30) (Fig. 36) (Fig. 42)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 13) (Fig. 19) (Fig. 25) (Fig. 31) (Fig. 37) (Fig. 43)
Near Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 14) (Fig. 20) (Fig. 26) (Fig. 32) (Fig. 38) (Fig. 44)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 15) (Fig. 21) (Fig. 27) (Fig. 33) (Fig. 39) (Fig. 45)
Table 1: Summary of results for the Two Stream equilibrium.
Multiple conclusions can be drawn.
  • Local solver with near-optimal initialization (good initial guess) perform universally well, and the stability is achieved in both under-parameter regime, and over-parameter regime, as seen in Figures 15, 21, 27, 33, 39 and 45. This signifies the importance of making a good initial guess. Making a good initial guess is usually not a feasible task for a classical PDE-constrained optimization, but is possible in this context by deploying the study of the dispersion relation, see discussion in Section 3.
  • Adaptive solver is not compatible with the objective functions (EE) and (EET), as seen in Figures 22, 40, 18, 24, 36, 42, 38 and 44. This is already suggested by Figure 6, (EET) where non-physical local minima can be found when parameters take on very large values. When adaptive solvers such as GD with line-search are deployed, the searching can take the iteration very far into these large parameter regions. The found external field is extremely strong and it completely dominates the whole plasma behavior. When it happens, the self-generated electric field is indeed very small, but the solution renders meaningless physics. The only exception in this case is to set the initial data from a near-optimal initialization. The basin is convex enough to control the line-search’s performance, see Figures 26 and 20.
  • The KL divergence does not seem to be a good metric for suppressing instability. The solution with good suppression is only observed when initialization is near-optimal already (see Figures 1415 and 33). With neither adaptive or local solver, in neither under-parameterized or over-parameterized regime, KL metric return a satisfactory solution as long as initialization is not near-optimal.
  • Counter-intuitively, over-parametrization in this setting does not seem to bring too much benefit (comparing the last three columns with the first three columns), with the only exception being the use of an adaptive solver for minimizing KL with a far-away initialization (comparing Figure 28 with Figure 10).
  • Without good initial guess, the only good solution is found by optimizing (EET) with simple GD solver in the properly parametrized setting when initialization is prepared in the mid-range (see Figure 25).
5.2 Bump-on-Tail example
For the Bump-on-Tail example, we conduct the same experiments and summarize them in Table 2 and Appendix B.2. Most observations drawn from the previous example still hold true. Yet we should identify the distinct features. In particular, our simulation seems to suggest that Bump-on-Tail instability is easier to be controlled, though the perfect recovery is more rare. More specifically, for Two Stream instability, when we have good initial guess, we obtained exact recovery when we use adaptive solvers in the under-parameterized regime and local solvers in the over-parameterized regime. We fail to do so in the Bump-on-Tail situation except for (EET). On the other hand, we find many more situations in which, though we are not being able to find exact recovery, the solutions are still relatively stable. This strongly suggests the local minima are good controls nevertheless, see trajectory data in Figures 46, 47, 59, 53, 51, 63 and 57.
Finding good initialization is still instrumental and it brings out-performance across solvers and regimes. Adaptive solvers are still incompatible with (EE) and (EET), and the performance is universally bad. Over-parameterization still does not necessarily bring benefits. In particular, with good initial guess and local solver, looking for best control in the over-parameterized regime brings even worse reconstruction, see Figures 69 and 81 vs. Figures 51 and 63.
However, in comparison to the Two Stream case, KL divergence seems to be a reasonable objective function, by minimizing which, acceptable controls for Bump-on-Tail can still be found. This resonates our earlier discussion on local minima serving as good controls for Bump-on-Tail instability.
Under-parametrized (5.2) Over-parametrized (5.3)
Init. Step (KL) ($E$) (EET) (KL) ($E$) (EET)
Far Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 46) (Fig. 52) (Fig. 58) (Fig. 64) (Fig. 70) (Fig. 76)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 47) (Fig. 53) (Fig. 59) (Fig. 65) (Fig. 71) (Fig. 77)
Mid Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 48) (Fig. 54) (Fig. 60) (Fig. 66) (Fig. 72) (Fig. 78)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 49) (Fig. 55) (Fig. 61) (Fig. 67) (Fig. 73) (Fig. 79)
Near Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 50) (Fig. 56) (Fig. 62) (Fig. 68) (Fig. 74) (Fig. 80)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 51) (Fig. 57) (Fig. 63) (Fig. 69) (Fig. 75) (Fig. 81)
Table 2: Summary of results for the Bump-on-Tail equilibrium.
Acknowledgements
Q.L. and Y.Y. acknowledge support from NSF-DMS-2308440. M.G. acknowledges support from NSF-DMS-2012292.
Appendix A Comparison between $L^{2}$ and KL
Another possible function is to use the $L^{2}$ disparity:
$\mathcal{J}(f[H])=\frac{1}{2}\|f_{T}[H]-f_{\text{eq}}\|_{L^{2}(x,v)}^{2}=\frac{1}{2}\int_{-L_{v}}^{L_{v}}(f[H](T,x,v)-f_{\text{eq}}(v))^{2}\mathrm{d}x\mathrm{d}v$ ($L^{2}$)
$\mathcal{J}(f[H])=\frac{1}{2}\int_{0}^{T}\int_{-L_{v}}^{L_{v}}(f[H](t,x,v)-f_{\text{eq}}(v))^{2}\mathrm{d}x\mathrm{d}v\mathrm{d}t\,.$ ($L^{2}$T)
In our simulation, we find the numerical behavior of these objective functions are rather similar to those of KL ((KL) and (KLT)). We provide numerical evidence in this section.
Figure 9: Landscape of Two Stream instability for $b_{1}$ for (4.1)(top) and Bump-on-Tail instability for $a_{1}$ for (4.3)(bottom) on the domain $[0.07,0.07]$ of the objectives (KL)(left), ($L^{2}$)(center-left), (KLT)(center-right) and ($L^{2}$T)(right).
In Figure 9, we show the one-dimensional case for both examples. For the Two Stream case, both landscapes exhibit no apparent differences, for both integrated-in-time case, and the disparity computed from the last frame. The situation for the Bump-on-Tail case is similar, with landscape for ($L^{2}$) agreeing with that for (KL) and ($L^{2}$T) agreeing with (KLT). We should note, however, ($L^{2}$) seems to show deeper local minima, making the computation more difficult to find the global optimal solution.
Appendix B Summary of results
In this section we present in detail the results seen in Tables 1 and 2.
B.1 Two Stream example
Figure 10: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 11: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 12: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 13: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 14: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 15: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 16: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 17: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
Figure 18: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 19: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 20: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 21: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 22: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 23: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
Figure 24: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 25: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 26: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 27: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 28: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 29: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 30: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 31: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 32: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 33: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 34: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 35: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 36: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 37: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 38: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 39: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 40: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 41: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 42: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 43: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 44: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 45: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
B.2 Bump-on-Tail example
Figure 46: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 47: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 48: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 49: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 50: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 51: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 52: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 53: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 54: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 55: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 56: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$,$H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 57: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 58: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 59: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 60: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 61: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 62: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$,$H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 63: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 64: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 65: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 66: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 67: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 68: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 69: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 70: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
Figure 71: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 72: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 73: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 74: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 75: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 76: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
Figure 77: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 78: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 79: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 80: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 81: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).